Elaine is 8 years younger than her brother Tyler. The sum of their ages is 34.Define a variable and solve the equation to find the age of Elaine and Tyler
Answer by richwmiller(17219)

elaine=tyler+(-8)
elaine+0+tyler+0=34
tyler+(-8)+0+tyler+0=34
tyler=(34-(0)-(-8)/2
tyler=21
elaine=21+(-8)=13
tyler=21 elaine=13